I'm having a hard time with this problem. I'm not sure if I'm doing this right.
It says solve the system by the elimination method
7x+5y=42
4x-3y=24
doing this I came up with this 28x+20y=168
28x-21y=168
some how I end up with -1. I'm trying to cancell out the x but the 168 is bothering me, it doesnt look right. Please help me

Answer by 1980.vikas(7)

first of all i want to say in elimination method we have to eliminate x or y in any of the equation.
let us eliminate x from 4x - 3y= 24, we get
x = 6 + (3/4)y ....................(i)
now put this value of x in the other equation i.e. 7x + 5y = 42
=> 7 * (6 + (3/4)y)+ 5y = 42
=> 42 + (21/4)y + 5y = 42
=> (21/4)y + 5y = 0
=> y[(21/4) + 5] = 0
=> y = 0
Putting the value of y in the equation (i), we get:
x = 6